# Topic-Specific Classifiers are Better Relevance Judges than Prompted LLMs

### v1

@@ 5 Results & Discussion
5.1 Model Judges from Deep Relevance Judgments
In the first experiment, we investigate whether ranking-models can successfully be specialized as single-topic judges in a deeply judged setting, i.e., without restricting the quantity of available training data.
Can ranking models be finetunged to judge topic-specific relevance?
Using the train/test set as described in Section˜4, we first evaluate the fully finetuned topic judge models w.r.t. precision, recall, F1, and accuracy. Results are shown in the first row of Figure˜1. Overall, training topic-specific judge models is successful across all tested model types. Their effectiveness exceeds 0.88 average F1 in all cases, with near-perfect recall as well as accuracy. Lower precision reflects class imbalance: only 6% of Robust04 documents are relevant, causing false positive bias despite weighted loss training. Between the different model types, mono-decoders perform best, exhibiting both the highest average scores, as well as the lowest variance over topics. Cross-encoders are second, scoring slightly higher than bi-encoders on average, yet also exhibiting greater variance.
Figure 1: Effectiveness of our topic-specific ranker approach in terms of F1, precision, recall, and accuracy when fine-tuning versus when using LoRA between different model architectures.
@@ 5.1 Model Judges from Deep Relevance Judgments
To reduce the computational overhead in training judge models, and making them ‘portable’ for dissemination, we investigate the use of LoRA adapters instead of full finetuning. Using the same evaluation procedure as before, the second row of Figure˜1 shows that adapter-based judges only exhibit minimal deviation in evaluation metrics compared to their fully finetuned counterparts. Accuracy and recall remain the same, while a slight drop of 0.05 - 0.1 in precision (and thus F1) is observable, varying by model type. Cross-encoder adapters perform worse than bi-encoder and mono-decoder adapters, both in absolute terms and in relative difference to full finetuning. Mono-decoder adapters remain the best choice overall, and their adapter versions prove to be a suitable replacement for full finetuning.
Figure 2: Precision, recall, F1, and Accuracy for binary relevance classification by adapter models using different amounts of training samples. Colors indicate datasets: X TREC Robust04, X TREC DL19, X TREC DL20. DL dataset are in-domain (MSMARCO) for base ranking model, Robust04 is not.
Figure 3: Spearmans’ $\rho$ rank correlation of system orderings at different depths of nDCG to human judgments at different run-subsampling rates for simulated pooling. Missing relevance judgments infilled by X assuming non-relevance (baseline, dashed), and judge adapter relevancy predictions for X $t=256$, X $t=192$, X $t=128$, X $t=64$, where $t$ is adapter training set size. Bootstrapped uncertainty as shaded areas ($n=20$).
5.2 Model Judges from Shallow Relevance Judgments
@@
We compare the four variants of judge adapters, trained with 64, 128, 192, or 256 training examples, respectively, by their accuracy, precision, recall, and F1 on the remaining judgment data per topic. Figure˜2 shows that results are similar across all metrics and models for DL19 and DL20 datasets, reaching near-perfect precision at slightly lower but still excellent recall and accuracy. Recall on Robust04 remains comparable to the DL datasets, yet accuracy is reduced in low-data variants, and precision drops to between 0.7 and 0.75 on average. Overall, we see a trend of more training data yielding better predictions at diminishing returns for >128>128 samples. The high scores throughout assert the viability of judge adapters even on shallow pools.
@@
To measure the impact the use of judge adapters has on system rankings in a realistic scenario, we simulate missing judgments through reduced pools, and then compare the rank correlation between system rankings derived from predicted judgments to those derived from the ground-truth human relevance judgments. Reduced pools are simulated through run subsampling, where the top 100 documents of all chosen runs are assigned their ground-truth scores, i.e., the train set for the adapter model, while the additional documents from non-chosen runs are treated as ‘new’, i.e., unjudged, and receive scores as predicted by the respective trained topic judge adapter. This allows us to simulate realistic unjudged documents patterns. In a bootstrapping-like setup, each such simulation is repeated 20 times with different seeds. Note that very small pools (1-3 runs) might not include enough documents to train some adapters with larger train set sizes, so these are omitted. As baseline and lower bound to compare to, we use evaluation results where missing relevance judgments are deemed irrelevant by default.
Figure˜3 plots the attained correlation scores over run subsampling rates for the different datasets and adapter train set sizes. Throughout, the system rankings on predicted relevance judgments very strongly correlate with the ground truth system ranking. Here, too, higher amounts of training data yield more reliable judge adapters. As pooling sizes increases, i.e., the ratio of predicted to ground-truth relevance labels shifts towards the ground-truth, the ranking quickly approach perfect correlation (>0.99>0.99) at run subsampling rates of 0.2 and above across all datasets. However, also with extremely sparse pools of around 3-4 runs, given enough training data in each topic, judge adapters can successfully infer the correct system ranking, reaching correlation values of around 0.95. In comparsion, the naïve common practice of deeming unjudged documents as non-relevant scores worse for extremely shallow pools, while difference even out as the proportion of ground-truth judgments increases. Judge adapters furthermore offer much stabler results, showing less variance across the 20 iterations of run subsampling.
@@
Comparing the effectiveness of judges across the three datasets as per Figure˜2, we can observe differences coinciding with data domain: judges are consistently better in the two in-domain evaluation datasets DL19 and DL20, which both derive from MSMARCO [15], the corpus the monoT5 base ranker was trained on, opposed to Robust04, which is more out-of-domain. The same trend is apparent in Figure˜3, where both the absolute attained scores are lower for Robust04, and the relative difference to the baseline is smaller. Yet, while using effective in-domain base models is favorable for judge training, transferring rankers to other data domains is still feasible, but might require slightly more initial human judgments.
@@
Table 2: Spearmans $\rho$ rank correlation between system rankings from original judgments and predicted relevance, for different depths of nDCG; and Krippendorff’s $\alpha$ agreement for nominal scales between original judgments and predicted relevance.
Approach Correlation Agreement
nDCG@5 nDCG@10 nDCG@50
monoT5 Adapter ($t=64$) 0.843 7760.843\,776 0.813 6760.813\,676 0.871 7120.871\,712 0.552 9010.552\,901
monoT5 Adapter ($t=128$) 0.869 9010.869\,901 0.884 3370.884\,337 0.876 0960.876\,096 0.489 9650.489\,965
monoT5 Adapter ($t=192$) 0.943 1330.943\,133 0.879 0770.879\,077 0.899 1230.899\,123 0.559 8730.559\,873
monoT5 Adapter ($t=256$) 0.966 5690.966\,569 0.954 2370.954\,237 0.952 4250.952\,425 0.688 4270.688\,427
Zero-Shot GPT-4 [2] 0.309 4680.309\,468 0.095 8500.095\,850 0.741 4960.741\,496 0.154 4090.154\,409
Zero-Shot GPT-4 [28] 0.632 9050.632\,905 0.602 5130.602\,513 0.826 8260.826\,826 0.457 1460.457\,146
Few-Shot GPT-4 ($n=8$, [2]) 0.345 4700.345\,470 0.206 4290.206\,429 0.765 9850.765\,985 0.200 4170.200\,417
Few-Shot GPT-4 ($n=8$, [28]) 0.803 2140.803\,214 0.739 1000.739\,100 0.886 1480.886\,148 0.489 7210.489\,721
Baseline (0-filling) 0.183 928-0.183\,928 0.279 953-0.279\,953 0.521 3910.521\,391 0.476 321-0.476\,321
5.3 Comparison to LLM-as-a-judge
@@
In the third experiment, we compare our proposed approach to establish LLM-as-a-judge methods, which have recently been proposed as effective solution to the unjudged document problem. Furthermore, given the demonstrated benefits of topic specialization for judges, we investigate if establishing specific topic context through few-shot prompting increases LLM-as-judge effectiveness.
@@
We sample 256 documents per topic as train set for all topics of DL20. These retain their ground truth relevance annotations, and are used both to train judge adapters, as well as sample 4 relevant and 4 non-relevant few-shot examples from. In a similar evaluation setup as in the previous experiment, we compare Spearmans’ $\rho$ rank correlation between system rankings as induced by different depths of nDCG. Additionally, we calculate the overall agreement between each approach and the ground truth labels using Krippendorff’s $\alpha$ for nominal scales. Table˜2 lists results. Judge adapters provide the best approximation of the original pool, exhibiting both highest correlation (>0.81>0.81) and agreement (>0.55>0.55) values in all cases. With the exception of nDCG@50 for lower train set sizes, across all metrics, judge adapters outscore the LLM-as-a-judge approaches. Furthermore, judge adapter provide more stable correlation values throughout the different evaluation depths.
Does topic context through few-shot prompting increase LLM-as-judge effectiveness?
Comparing zero-shot and few-shot variants of LLM-as-a-judge in Table˜2, in both cases, providing examples yields increased effectiveness. This provides further evidence for our underlying assumption that topic specialization is a favorable characteristic. Yet, few-shot prompting massively increases the token budget and thus computational cost needed for relevance label inference, while judge adapters have negligible inference cost in comparison.
6 Practical Considerations for Using Judge Adapters

### v2

@@ 5. Results & Discussion
5.1. Model Judges from Deep Relevance Judgments
In the first experiment, we investigate whether ranking models can successfully be specialized as single-topic judges in a deeply judged setting, i.e., without restricting the quantity of training data.
Can ranking models be finetuned to judge topic-specific relevance?
Using the train/test set as described in Section˜4, we first evaluate the fully finetuned topic judge models w.r.t. precision, recall, F1, and accuracy. Results are shown in the first row of Figure˜2. Overall, training topic-specific judge models is successful across all tested model types. Their effectiveness exceeds 0.88 average F1 in all cases, with near-perfect recall and accuracy. Lower precision reflects class imbalance: only 6% of Robust04 documents are relevant, causing false positive bias despite weighted loss training. Between the different model types, the mono-decoder was the most effective, exhibiting both the highest average scores and the lowest variance over topics. Cross-encoders are second, scoring slightly higher than bi-encoders on average, yet also exhibiting greater variance.
@@ 5.1. Model Judges from Deep Relevance Judgments
To reduce the computational overhead in training judge models and make them “portable” for dissemination, we investigate the use of LoRA adapters instead of full finetuning. Using the same evaluation procedure as before, the second row of Figure˜2 shows that adapter-based judges only exhibit minimal deviation in evaluation metrics compared to their fully finetuned counterparts. Accuracy and recall remain the same, while a slight drop of 0.05 - 0.1 in precision (and thus F1) is observable, varying by model type. Cross-encoder adapters perform worse than bi-encoder and mono-decoder adapters, both in absolute terms and in relative difference to full finetuning. Mono-decoder adapters remain the best choice overall, and their adapter versions are a suitable replacement for full finetuning.
5.2. Model Judges from Shallow Relevance Judgments
@@
We compare the four variants of judge adapters, trained with 64, 128, 192, or 256 training examples, respectively, by their accuracy, precision, recall, and F1 on the remaining judgment data per topic. Figure˜3 shows that results are similar across all metrics and models for DL19 and DL20 datasets, reaching near-perfect precision at slightly lower but still excellent recall and accuracy. Recall on Robust04 remains comparable to the DL datasets, yet accuracy is reduced in low-data variants, and precision drops to between 0.7 and 0.75 on average. Overall, we see a trend of more training data yielding better predictions at diminishing returns for >128>128 samples. The high scores throughout assert the viability of judge adapters even on shallow pools.
Figure 3. Precision, recall, F1, and Accuracy for binary relevance classification by adapter models using different amounts of training samples. Colors indicate datasets: X TREC Robust04, X TREC DL19, X TREC DL20. The DL datasets are in-domain (MSMARCO) for the base ranking model, Robust04 is not.
@@
To measure the impact using judge adapters has on system rankings in a realistic scenario, we simulate missing judgments through reduced pools, and then compare the rank correlation between system rankings derived from predicted judgments to those derived from the ground-truth human relevance judgments. Reduced pools are simulated through run subsampling, where the top 100 documents of all chosen runs are assigned their ground-truth scores, i.e., the train set for the adapter model, while the additional documents from non-chosen runs are treated as ‘new’, i.e., unjudged, and receive scores as predicted by the respective trained topic judge adapter. This allows us to simulate realistic unjudged documents patterns. In a bootstrapping-like setup, each such simulation is repeated 20 times with different seeds. Note that very small pools (1-3 runs) might not include enough documents to train some adapters with larger train set sizes, so these are omitted. As baseline and lower bound to compare to, we use evaluation results where missing relevance judgments are deemed irrelevant by default.
Figure 4. Spearmans’ $\rho$ rank correlation of system orderings at different depths of nDCG to human judgments at different run-subsampling rates for simulated pooling. Missing relevance judgments infilled by X assuming non-relevance (baseline, dashed), and judge adapter relevancy predictions for X $t=256$, X $t=192$, X $t=128$, X $t=64$, where $t$ is adapter training set size. Bootstrapped uncertainty as shaded areas ($n=20$).
Figure˜4 plots the correlation values over run subsampling rates for the different datasets and adapter train set sizes. Throughout, the system rankings on predicted relevance judgments very strongly correlate with the ground truth system ranking. Here, too, higher amounts of training data yield more reliable judge adapters. As pooling sizes increases, i.e., the ratio of predicted to ground-truth relevance labels shifts towards the ground-truth, the ranking quickly approach perfect correlation (>0.99>0.99) at run subsampling rates of 0.2 and above across all datasets. However, also with extremely sparse pools of around 3-4 runs, given enough training data in each topic, judge adapters can successfully infer the correct system ranking, reaching correlation values of around 0.95. In comparison, the naïve common practice of deeming unjudged documents as non-relevant scores worse for extremely shallow pools, while difference even out as the proportion of ground-truth judgments increases. Judge adapters furthermore offer much stabler results, showing less variance across the 20 iterations of run subsampling.
@@
Comparing the effectiveness of judges across the three datasets as per Figure˜3, we can observe differences coinciding with data domain: judges are consistently better in the two in-domain evaluation datasets DL19 and DL20, which both derive from MSMARCO (nguyen2016msmarco), the corpus the monoT5 base ranker was trained on, opposed to Robust04, which is more out-of-domain. The same trend is apparent in Figure˜4, where both the absolute attained scores are lower for Robust04, and the relative difference to the baseline is smaller. Yet, while using effective in-domain base models is favorable for judge training, transferring rankers to other data domains is still feasible, but might require slightly more initial human judgments.
@@
Table 3. Comparison between original human judgments and and relevance predicted by each automatic approach (A = archetype, P = prompt, M = model, R = reasoning). Spearmans $\rho$ rank correlation between system rankings from original judgments and predicted relevance per topic, mean and 95% CI over topics, for different depths $k$ of different retrieval measures; and Krippendorff’s $\alpha$ agreement for nominal scales between original judgments and predicted relevance. Highest per archetype and column marked bold. Highest adapter colored green; highest LLM-as-a-judge colored blue.
Approach $\rho$ / NDCG@k $\rho$ / Prec@k $\rho$ / Rec@k $\rho$ / RRank@k $\alpha$
A P M R $k=10$ $k=50$ $k=10$ $k=50$ $k=10$ $k=50$ $k=10$ $k=50$
Baseline (0-filling) $0.33\pm 0.17$ $0.49\pm 0.13$ $0.52\pm 0.10$ $0.70\pm 0.09$ $0.52\pm 0.10$ $0.70\pm 0.09$ $0.52\pm 0.19$ $0.44\pm 0.14$ 0.335 331-0.335\,331
Zero-Shot LLM (arabzadeh:2025) Ministral-3 $0.76\pm 0.07$ $0.85\pm 0.04$ $0.84\pm 0.04$ $0.87\pm 0.04$ $0.84\pm 0.04$ $0.87\pm 0.04$ $0.76\pm 0.08$ $0.79\pm 0.07$ 0.485 5820.485\,582
(arabzadeh:2025) Qwen3-Next $0.81\pm 0.05$ $0.86\pm 0.05$ $6\pm 0.04 0.88±0$$8\pm 0.04$ $6\pm 0.04 0.88±0$$8\pm 0.04$ $0.82\pm 0.07$ $0.83\pm 0.06$ 0.546 0540.546\,054
(arabzadeh:2025) GPT-OSS $0.79\pm 0.07$ $0.84\pm 0.05$ $4\pm 0.05 0.88±0$$8\pm 0.05$ $4\pm 0.05 0.88±0$$8\pm 0.05$ $0.83\pm 0.08$ $0.84\pm 0.07$ 0.497 9430.497\,943
(arabzadeh:2025) Minimax-M1 $0.77\pm 0.07$ $0.83\pm 0.05$ $0.84\pm 0.05$ $0.87\pm 0.05$ $0.84\pm 0.05$ $0.87\pm 0.05$ $0.78\pm 0.09$ $0.80\pm 0.08$ 0.485 0850.485\,085
(arabzadeh:2025) GPT-4o $0.73\pm 0.08$ $0.82\pm 0.06$ $0.79\pm 0.06$ $0.85\pm 0.06$ $0.79\pm 0.06$ $0.85\pm 0.06$ $0.81\pm 0.08$ $0.82\pm 0.07$ 0.395 0210.395\,021
(upadhyay2024umbrela) Ministral-3 $0.82\pm 0.05$ $0.84\pm 0.06$ $0.87\pm 0.03$ $0.87\pm 0.05$ $0.87\pm 0.03$ $0.87\pm 0.05$ $0.81\pm 0.07$ $0.81\pm 0.07$ 0.332 3890.332\,389
(upadhyay2024umbrela) Qwen3-Next $0.82\pm 0.05$ $0.85\pm 0.05$ $7\pm 0.03 0.88±0$$8\pm 0.04$ $7\pm 0.03 0.88±0$$8\pm 0.04$ $0.81\pm 0.06$ $0.82\pm 0.06$ 0.533 2970.533\,297
(upadhyay2024umbrela) GPT-OSS $0.82\pm 0.05$ $5\pm 0.06 0.88±0$$8\pm 0.04$ $8\pm 0.05 0.88±0$$8\pm 0.04$ $0.88\pm 0.05$ $0.81\pm 0.07$ $0.81\pm 0.07$ 0.490 3450.490\,345
(upadhyay2024umbrela) Minimax-M1 $0.84\pm 0.05$ $0.84\pm 0.07$ $9\pm 0.03 0.88±0$$8\pm 0.05$ $9\pm 0.03 0.88±0$$8\pm 0.05$ $0.83\pm 0.07$ $0.84\pm 0.06$ 0.462 8120.462\,812
[…]
Few-Shot LLM ($n=8$) (arabzadeh:2025) Ministral-3 $0.64\pm 0.08$ $0.79\pm 0.06$ $0.75\pm 0.06$ $0.86\pm 0.04$ $0.75\pm 0.06$ $0.86\pm 0.04$ $0.60\pm 0.10$ $0.61\pm 0.10$ 0.319 7310.319\,731
[…]
(arabzadeh:2025) Minimax-M1 $0.79\pm 0.07$ $0.84\pm 0.06$ $0.85\pm 0.05$ $0.87\pm 0.05$ $0.85\pm 0.05$ $0.87\pm 0.05$ $0.80\pm 0.09$ $0.81\pm 0.08$ 0.514 4330.514\,433
(arabzadeh:2025) GPT-4o $0.73\pm 0.09$ $0.82\pm 0.07$ $0.79\pm 0.07$ $0.86\pm 0.06$ $0.79\pm 0.07$ $0.86\pm 0.06$ $0.81\pm 0.09$ $0.82\pm 0.08$ 0.425 4770.425\,477
(upadhyay2024umbrela) Ministral-3 $0.81\pm 0.05$ $0.87\pm 0.04$ $0.87\pm 0.04$ $0.89\pm 0.04$ $0.87\pm 0.04$ $0.89\pm 0.04$ $0.81\pm 0.07$ $0.81\pm 0.06$ 0.471 2750.471\,275
[…]
(upadhyay2024umbrela) GPT-OSS $0.85\pm 0.05$ $0.83\pm 0.08$ $0.89\pm 0.03$ $0.87\pm 0.06$ $0.89\pm 0.03$ $0.87\pm 0.06$ $0.81\pm 0.07$ $0.81\pm 0.07$ 0.489 2230.489\,223
(upadhyay2024umbrela) Minimax-M1 $0.83\pm 0.05$ $6\pm 0.06 0.88±0$$8\pm 0.03$ $8\pm 0.04 0.88±0$$8\pm 0.03$ $0.88\pm 0.04$ $0.83\pm 0.07$ $0.83\pm 0.06$ 0.468 2090.468\,209
(upadhyay2024umbrela) GPT-4o $0.83\pm 0.05$ $7\pm 0.05 0.88±0$$8\pm 0.04$ $9\pm 0.04 0.88±0$$8\pm 0.04$ $0.89\pm 0.04$ $0.81\pm 0.08$ $0.83\pm 0.07$ 0.507 9970.507\,997
[…]
monoT5 ($t=128$) $0.84\pm 0.05$ $0.96\pm 0.01$ $0.89\pm 0.04$ $0.98\pm 0.01$ $0.89\pm 0.04$ $0.98\pm 0.01$ $0.86\pm 0.06$ $0.87\pm 0.06$ 0.807 5800.807\,580
monoT5 ($t=192$) $0.85\pm 0.05$ $0.96\pm 0.02$ $0.90\pm 0.04$ $0.98\pm 0.01$ $0.90\pm 0.04$ $0.98\pm 0.01$ $0.89\pm 0.06$ $0.90\pm 0.06$ 0.830 5810.830\,581
[…]
5.3. Comparison to LLM-as-a-judge
@@
In the third experiment, we compare our proposed approach to established LLM-as-a-judge methods, which have recently been proposed as effective solutions to the unjudged document problem. Furthermore, given the demonstrated benefits of topic specialization for judges, we investigate whether establishing specific topic context through few-shot prompting increases LLM-as-a-judge effectiveness. As before, we sample 256 documents per topic as train set for all topics of DL20. These retain their ground truth relevance annotations, and are used both to train judge adapters, as well as to sample 4 relevant and 4 non-relevant few-shot examples from. We then use each of the five different LLMs, each with both prompt variants and in zero- and few-shot configurations, to judge the remaining unjudged documents from the TREC DL20 pool.
To evaluate label source quality, we compare system rankings induced by different label sources at varying retrieval depths ($k\in\{5,10,50\}$) across five metrics: Reciprocal Rank (RR), normalized Discounted Cumulative Gain (NDCG), Mean Average Precision (MAP), Recall, and Precision. The systems being ranked are the original TREC DL20 submissions; only the relevance judgments differ across label sources. For each metric, we calculate Spearman’s $\rho$ rank correlation between system rankings computed using complete human label ground truth versus system rankings computed using the hybrid judgment sets (initial 256 human judgments plus automatic labels for remaining documents). Correlation is measured per-topic and then averaged over all evaluation topics of TREC DL20. Additionally, we measure direct inter-annotator agreement between each automatic labeling approach and the ground truth labels using Krippendorff’s $\alpha$ for nominal scales. Agreement is measured on the complete label set across all topics. Table˜3 lists results. We interpret our findings through the lens of LLM evaluation tropes proposed by (dietz2025principles), which identify common failure modes when using LLMs as evaluators.
@@
Judge adapters provide the best approximation of human judgments, exhibiting both highest rank correlations and inter-annotator agreement ($\alpha=0.81$$0.88$) across all configurations. At $k=10$, adapters achieve $\rho\geq 0.84$ across all metrics; at $k=50$, correlations reach $\rho\geq 0.96$ for NDCG, Precision, and Recall. RR@k correlations are notably lower ($\rho=0.87$$0.93$). Across all metrics and evaluation depths, judge adapters consistently outscore LLM-as-a-judge approaches for completing the judgment pool.
While differences in system rankings between judge adapters and LLM-as-a-judge remain moderate ($\Delta_{k=10}\approx 0.04,\Delta_{k=50}\approx 0.09$), the inter-annotator agreement for LLMs is much lower ($\alpha=0.32$$0.58$) compared to adapters. Generally, $\alpha\geq 0.80$ indicates good reliability, while $0.67\leq\alpha<0.80$ allows only tentative conclusions (artstein2008intercoder). This exemplifies what (dietz2025principles) term the Ignored Label Correlation trope: high system-level correlation can obscure important disagreements at the judgment level (faggioli2023perspectives). LLMs seemingly apply a consistent but more lenient relevance threshold, flagging substantially more documents as relevant than human annotators, but do so systematically across all systems rather than favoring particular retrieval approaches. Thus, LLMs preserve the relative ordering of systems. Examining correlation patterns across depths $k\in\{10,50\}$, this further indicates that with deeper unjudged pools, LLMs may introduce more false positive ratings of relevance, thus producing lower system-level correlation compared to the adapter approaches. We further observe that RR@k exhibits the largest spread across label sources, particularly when few-shot prompting destabilizes smaller models (e.g., Ministral-3 drops to $\rho=0.60$ for RR@10). NDCG@k, Recall@k, and Precision@k remain comparatively stable. This suggests that disagreements between label sources primarily manifest in the very top ranks, affecting which systems appear best at identifying the single most relevant document, while set-level assessment remains robust.
Furthermore, the naive baseline of assuming non-relevance, which exhibits negative agreement ($\alpha=-0.34$), achieves passable system-level correlation at $k=50$ (Precision, Recall: $\rho=0.70$). This provides further evidence that system-level orderings are rather robust while on the label-level, judgments can be drastically different, corroborating the interpretation that $\alpha$-values are more indicative of label source quality.
Does the backbone LLM model influence LLM-as-a-judge effectiveness?
Model choice demonstrates measurable but inconsistent effects on both ranking correlation and inter-annotator agreement. No clear pattern emerges favoring larger or reasoning-capable models: Qwen3-Next, a smaller open-weight model without reasoning, obtains the highest $\alpha$ values (0.530.530.580.58), while GPT-4o, identified as best annotation model by (arabzadeh:2025) shows surprisingly variable agreement ($\alpha=0.40$$0.54$) depending on prompt choice, even at its much higher parameter count and reasoning capability. This suggests that relevance assessment may be a calibration task rather than a reasoning task. Inter-annotator agreement ranges from $\alpha=0.33$ to $0.55$ across zero-shot configurations. This pattern suggests that while different models disagree on the threshold for relevance, they maintain similar system ranking behavior, i.e., do not favor particular documents or contributing systems, which would manifest in diverging system orderings.
However, this similarity may indicate susceptibility to the Loss of Variety of Opinion trope: if all LLMs converge on similar ranking behaviors due to shared training paradigms or data sources, the community loses the diversity of perspectives that human assessors provide, and which topic-specific adapters can reproduce. Furthermore, while the training data for our adapter backbone model is known, we cannot guarantee that the open-weight LLMs did not see test set data from TREC-DL20 during pretraining, i.e., a threat of Test Set Leak as identified by (dietz2025principles), inflating their scoring effectiveness on this collection.
Does prompt variation influence LLM-as-a-judge effectiveness?
Comparing the two prompt variants across identical model configurations, we observe minimal systematic differences in ranking correlations. Both the prompt of (arabzadeh:2025) and (upadhyay2024umbrela) achieve comparable $\rho$ values (typically within 0.010.010.030.03), with neither consistently outperforming the other. This stability suggests that system ranking validity is relatively robust to prompt formulation. However, inter-annotator agreement varies substantially ($\Delta\alpha$ up to $0.15$ for identical models with different prompts, e.g., Ministral-3), indicating that prompt design affects the relevance threshold applied by the LLM. Comparing zero-shot and few-shot variants, few-shot prompting yields mixed results. For Qwen3-Next with (arabzadeh:2025)’s ((arabzadeh:2025)) prompt, few-shot examples increase agreement from $\alpha=0.55$ to $\alpha=0.58$. However, for Ministral-3 with the same prompt, few-shot prompting causes substantial degradation (RR@10 drops from $0.76$ to $0.60$; $\alpha$ drops from $0.49$ to $0.32$). Overall, few-shot prompting at $n=8$ examples does not yield consistent improvements. At the same time, it increases token budget and computational cost. The susceptibility of $\alpha$-values to model and prompt choice illustrates (dietz2025principles)’s ((dietz2025principles)) LLM Evolution trope, i.e., evaluation outcomes depend on model choice, configuration, and prompt usage.
Statistical Significance
Table˜4 lists significance levels for comparing adapters against the best-performing LLM per archetype-prompt configuration. These confirm that adapters significantly outperform LLM-as-a-judge at greater evaluation depths ($k=50$, $p<0.001$ in all cases), while with shallow depth ($k=10$) differences are statistically significant for 3 out of the 4 comparisons for $t=256$, and 2 out of 4 for $t=64$. Since the comparisons are not independent (sharing the same system pool), and we test against the best-performing LLM rather than all configurations (a conservative choice), we do not apply multiple testing correction. However, even under Bonferroni correction, the $k=50$ results remain highly significant ($0.05/16=0.003$). This reinforces that adapters are reliably advantageous to LLM-as-a-judge when evaluating beyond the very top ranks, i.e., extending (shallow) human-judged pools.
Table 4. Significance level for comparing system correlations between adapter judges and best-performing LLM-as-a-judge approach per archetype and prompt for NDCG@10 and NDCG@50; paired one-sided $t$-test, $n=54$. Non-significant values ($p>0.05$) greyed out.
Adapter (Ours)
$t=64$ $t=128$ $t=192$ $t=256$
LLM-as-a-judge NDCG@10 Qwen3-Next, (arabzadeh:2025) 0.044 0210.044\,021 0.193 7850.193\,785 0.124 7800.124\,780 0.015 1020.015\,102
GPT-4o, (upadhyay2024umbrela) 0.107 0410.107\,041 0.322 9340.322\,934 0.215 8740.215\,874 0.023 8330.023\,833
Qwen3-Next, few-s., (arabzadeh:2025) 0.049 9420.049\,942 0.210 3490.210\,349 0.140 1600.140\,160 0.024 9020.024\,902
Qwen3-Next, few-s., (upadhyay2024umbrela) 0.309 6130.309\,613 0.607 4470.607\,447 0.493 0670.493\,067 0.124 2050.124\,205
NDCG@50 Qwen3-Next, (arabzadeh:2025) 0.000 0120.000\,012 0.000 0860.000\,086 0.000 0530.000\,053 0.000 0120.000\,012
GPT-4o, (upadhyay2024umbrela) 0.000 0450.000\,045 0.000 1540.000\,154 0.000 1060.000\,106 0.000 0220.000\,022
Qwen3-Next, few-s., (arabzadeh:2025) 0.000 0020.000\,002 0.000 0310.000\,031 0.000 0170.000\,017 0.000 0040.000\,004
Qwen3-Next, few-s., (upadhyay2024umbrela) 0.000 0760.000\,076 0.000 5080.000\,508 0.000 3840.000\,384 0.000 0740.000\,074
6. Practical Considerations for Using Judge Adapters